OpenStudy (anonymous):

what is the distributive property of -6(2U-1)

OpenStudy (mathstudent55):

In general, the distributive property is: a(b - c) = ab - ac Take the number or expression outside the parentheses and multiply by all the numbers or expressions inside the parentheses.

OpenStudy (mathstudent55):

-6(U - 1) = -6 * U - (-6) * 1

OpenStudy (anonymous):

im just looking for the answer is all cuzz i need help on a virtual test

OpenStudy (mathstudent55):

Multiply -6 and U. Multiply -1 and 1. Then subtract.

OpenStudy (anonymous):

thnx lol and do you know thanswer to this question Simplify the following expression. 2x − 6y + 3x2 + 7y − 14x

OpenStudy (goldphenoix):



OpenStudy (mathstudent55):

Combine like terms. Like terms have the same variable part. There is only one term with x^2, so just write it. Then combine 2x and - 14x Then combine -6y and 7y

OpenStudy (anonymous):

thnx that relly helped alot

OpenStudy (mathstudent55):

wlcm